The y come food , by U. C.S. com 200 51.00 torch by 2000, That is the many of food is and the many of clothing M e th
Answer #1

Lasperyes index represents the amount she has to spend in 2000 against 1990. If she consumes same amount of food and clothing as she did in 1990. Therefore, the lasperyes index for 2000 is given by

L = 100(Y')/Y

Where, Y' = represents the amount she spends in 2000 to consume same amount of food and clothing.

As provided in the question this relationship will hold true

PEF = Pcc

The budget constraint

PEF + PCC=Y

Price of food and clothing in 1990 was $ 1

S F +C = 1200

Therefore, C = F = 600

À X 001 = 1

X 001 = 1 3F +4.50 1200

L= 100 3 x 600+ 4.5 x 600 1200

3+4.5 = L = 100 x

L= 50 x 7.5

L= 375

Laspereyes cost of living index in 2000 = 375

Ideal index

I = 100(Y)/Y

Where, Y'' = 3F' + 4.5C'

The bundle (F',C') will be on the same indifference curve as (F,C) and the indifference curve at this point will be tangent to the budget line with slope -(Pf' , Pc'). These are the prices of year 2000. As we know she spends equal amount on food and cloth. Therefore, 3F' = 4.5C'.

C = F = 600

F'C' = 600 × 600

=FX -F = 360,000 4.5

F2 = 540,000

F = 540,000

F 734.85

Now calculating, C' = (3/4.5)×F'

C' = (3/4.5)×734.85 = 489.90

I = 100 × (3F' + 4.5C')/ 1200

I = 100 ×(3×734.85 + 4.5 × 489.9)/1200

I = 367.425

Ideal index, I = 367.43

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.